# Neural Concept Verifier: Scaling Prover-Verifier Games via Concept Encodings

### v1

@@ 1 Introduction
Concurrently, Concept Bottleneck Models (CBMs) have emerged as a powerful framework for interpretable machine learning, structuring predictions through intermediate interpretable concept encodings (Koh et al.,, 2020; Stammer et al.,, 2021). Despite their advantages, CBMs typically employ linear classifiers on top of concept encoding layers, thereby potentially restricting expressivity and failing on tasks requiring nonlinear interactions among concepts (e.g., XOR problems, counting or permutation invariance) (Mahinpei and et al.,, 2021; Kimura et al.,, 2024; Lee et al., 2019b, ).
@@ 2 Background
Prover-Verifier Games. PVGs were introduced by Anil et al., (2021) as a game-theoretic framework to encourage learning agents to produce testable justifications through interactions between an untrusted prover and a trusted verifier. Their work showed that under suitable game settings, the verifier can learn robust decision rules even when the prover actively attempts to persuade it of arbitrary outputs. Wäldchen et al., (2024) extended this idea to the Merlin-Arthur Classifier (MAC), which provides formal interpretability guarantees by bounding the mutual information between the selected features and the ground-truth label. Recently, Kirchner et al., (2024) applied a PVG-inspired approach to improve legibility of Large Language Model (LLM) outputs and Amit and et al., (2024) introduced self-proving models that leverage interactive proofs to formally verify the correctness of model outputs. PVG-style setups have also been explored in safety-focused learning protocols (Irving et al.,, 2018; Brown-Cohen et al.,, 2024; Głuch et al.,, 2024). These developments reflect a broader trend of utilising multi-agent learning (Pruthi et al.,, 2022; Schneider and Vlachos,, 2023; Du et al.,, 2023; Nair et al.,, 2023; Stammer et al., 2024a, ). Our work builds on this line of research by embedding PVGs into a concept-based classification framework, addressing the dimensionality bottleneck that limits PVGs in high-dimensional settings.
@@ 3.4 Verifiable Classification via the Merlin-Arthur Classifiers
The second core component of NCV is a verifiable classifier of the Merlin-Arthur (MA) setup (Wäldchen et al.,, 2024) originally inspired by Interactive Proof Systems (Goldwasser et al.,, 1989). This setup generally formalizes the idea of proving that a classification decision is supported by a sparse and informative set of features.
@@ 4.1.3 NCV Model Instantiations
NCB-based Neural Concept Verifier. As detailed in Section  3.3, for evaluations on CLEVR-Hans3 and CLEVR-Hans7, we utilize NCB (Stammer et al., 2024b, ) to extract unsupervised, discrete concepts. Specifically, we use NCB models that were pretrained on the CLEVR (Johnson et al.,, 2017) dataset. NCB’s resulting concept encodings are structured as a set of object-centric slots. To handle this structure, we employ a permutation-invariant Set Transformer (Lee et al., 2019a, ) as the verifier (Arthur). The provers (Merlin and Morgana) are implemented as independent neural networks, each parameterized by a Set Transformer. They receive the full concept slot matrix as input and output a sparse selection mask with 12 nonzero entries, indicating the active concept blocks provided to the verifier. All components are jointly trained using the Adam optimizer. Further details and ablation studies can be found in Suppl.  B.

### v2

@@ 1 Introduction
Concurrently, Concept Bottleneck Models (CBMs) have emerged as a powerful framework for interpretable machine learning, structuring predictions through intermediate interpretable concept encodings (Koh et al.,, 2020; Stammer et al.,, 2021). Despite their advantages, CBMs typically employ linear classifiers on top of concept encoding layers, thereby potentially restricting expressivity and failing on tasks requiring nonlinear interactions among concepts (e.g., XOR problems, counting or permutation invariance) (Mahinpei et al.,, 2021; Kimura et al.,, 2024; Lee et al.,, 2019).
@@ 2 Background
Prover-Verifier Games. PVGs were introduced by Anil et al., (2021) as a game-theoretic framework to encourage learning agents to produce testable justifications through interactions between an untrusted prover and a trusted verifier. Their work showed that under suitable game settings, the verifier can learn robust decision rules even when the prover actively attempts to persuade it of arbitrary outputs. Wäldchen et al., (2024) extended this idea to the Merlin-Arthur Classifier (MAC), which provides formal interpretability guarantees by bounding the mutual information between the selected features and the ground-truth label. Recently, Kirchner et al., (2024) applied a PVG-inspired approach to improve legibility of Large Language Model (LLM) outputs and Amit et al., (2024) introduced self-proving models that leverage interactive proofs to formally verify the correctness of model outputs. PVG-style setups have also been explored in safety-focused learning protocols (Irving et al.,, 2018; Brown-Cohen et al.,, 2024; Głuch et al.,, 2024). These developments reflect a broader trend of utilising multi-agent learning (Pruthi et al.,, 2022; Schneider and Vlachos,, 2024; Du et al.,, 2024; Nair et al.,, 2023; Stammer et al., 2024a, ). Our work builds on this line of research by embedding PVGs into a concept-based classification framework, addressing the dimensionality bottleneck that limits PVGs in high-dimensional settings.
@@ 3.4 Verifiable Classification via the Merlin-Arthur Classifiers
The second core component of NCV is a verifiable classifier of the Merlin-Arthur (MA) setup (Wäldchen et al.,, 2024) originally inspired by Interactive Proof Systems (Goldwasser et al.,, 1985). This setup generally formalizes the idea of proving that a classification decision is supported by a sparse and informative set of features.
@@ 4.1.3 NCV Model Instantiations
NCB-based Neural Concept Verifier. As detailed in Section  3.3, for evaluations on CLEVR-Hans3 and CLEVR-Hans7, we utilize NCB (Stammer et al., 2024b, ) to extract unsupervised, discrete concepts. Specifically, we use NCB models that were pretrained on the CLEVR (Johnson et al.,, 2017) dataset. NCB’s resulting concept encodings are structured as a set of object-centric slots. To handle this structure, we employ a permutation-invariant Set Transformer (Lee et al.,, 2019) as the verifier (Arthur). The provers (Merlin and Morgana) are implemented as independent neural networks, each parameterized by a Set Transformer. They receive the full concept slot matrix as input and output a sparse selection mask with 12 nonzero entries, indicating the active concept blocks provided to the verifier. All components are jointly trained using the Adam optimizer. Further details and ablation studies can be found in Suppl.  B.